Depreciation by Three Methods; Partial Years

Perdue Company purchased equipment on April 1 for $42,930. The equipment was expected to have a useful life of three years, or 5,940 operating hours, and a residual value of $1,350. The equipment was used for 1,100 hours during Year 1, 2,100 hours in Year 2, 1,800 hours in Year 3, and 940 hours in Year 4.

Required:

Determine the amount of depreciation expense for the years ended December 31, Year 1, Year 2, Year 3, and Year 4, by (a) the straight-line method, (b) units-of-output method, and (c) the double-declining-balance method.

Note: FOR DECLINING BALANCE ONLY, round the multiplier to four decimal places. Then round the answer for each year to the nearest whole dollar.
Answer #1

Calculate depreciation expense

Straight line Unit of output Double decline
Year 1 (42930-1350/3)*9/12 = 10395 1100*7 = 7700 42930*2/3*9/12 = 21465
Year 2 (42930-1350/12) = 13860 14700 7155+42930*1/3*2/3*9/12 = 14310
Year 3 13860 12600 2385+42930*1/3*1/3*2/3*9/12 = 4770
Year 4 3465 6580 1035
Total 41580 41580 41580